Suppose that insurance companies did a survey. They randomly surveyed 430 drivers and found that 330 claimed they always buckle up. We are interested in the population proportion of drivers who claim they always buckle up.

a) x = ? b) n = ? c) p' = ? d)

Which distribution should you use for this problem? (Round your answer to four decimal places.)

P- (_) (_,_)

e) Construct a 95% confidence interval for the population proportion who claim they always buckle up.

i. state the confidence interval

ii. sketch graph

iii. calculate error bound
Answer #1

Solution:

Given:

Sample size =n = 430

x = Number of drivers always buckle up = 330

Part a) x = ?

x = 330

Part b) n = ?

n = 430

Part c) p' = ?

\hat{p}=\frac{x}{n}

330 430

P=0.7674

Part d) Which distribution should you use for this problem?

Sampling distribution of sample proportions is approximate Normal distribution with mean of sample proportion:

Mộ=P

Mp= 0.7674


and standard deviation of sample proportions is:

px (1-P) Og = V!

0.7674 x (1 – 0.7674) Op = V 430

Op = 0.7674 X 0.2326 430

0 = V0.000415058

Op = 0.020373

Op = 0.0204

Thus

P ~ ( approximate Normal ) ( Mp= 0.7674 , Op = 0.0204 )

Part e) Construct a 95% confidence interval for the population proportion who claim they always buckle up.

-E, P+E)

where

E = Zcx op

We need to find zc value for c=95% confidence level.

Find Area = ( 1 + c ) / 2 = ( 1 + 0.95) /2 = 1.95 / 2 = 0.9750

Look in z table for Area = 0.9750 or its closest area and find z value.

0.1 V123 .00 .01 .02 .03 .04 .05 .06 0.0 ,5000 15040 5080 5120 .5160 5199 39 1.5398 .5438 5478 1.5517 .5557 .5596 .636 0.2 15

Area = 0.9750 corresponds to 1.9 and 0.06 , thus z critical value = 1.96

That is : Zc = 1.96

Thus

E = Zcx op

E = 1.96 x 0.020373

E = 0,0399

Thus

-E, P+E)

(0.7674-0.0399.0.7674 +0.0399)

(0.7275,0.8073)

i) a 95% confidence interval for the population proportion who claim they always buckle up is between (0.7275,0.8073)

ii) sketch graph

95% confidence interval P-E = 0.7275 = 0.7674 + E = 0.8073

iii) calculate error bound

E = Zcx op

E = 1.96 x 0.020373

E = 0,0399